OpenStudy (anonymous):

The area of the rectangle shown is 180 square cm. Find its dimensions. Use the equation: L(L + 3) = 180. (You may want to distribute first.) A rectangle with a length of L and a width of L + 3. A. 18 cm x 10 cm B. 12 cm x 15 cm C. 9 cm x 20 cm D. 6 cm x 30 cm

OpenStudy (anonymous):

L^2 + 3L = 180 L^2 + 3L - 180 = 0 (L - 12) ( L +15) = 0 L = 12 Answer is B

OpenStudy (phi):

Another way to do this: We assume that one of the choices is the correct answer. You know that one side is L and the other is L+3 (3 more than L). So look for a pair of numbers that differ by exactly 3. There is only one pair, choice B. You can then multiply the two numbers to see if it's 180. (Of course, if it isn't 180, then none of the answers work!)